# Disjoint Partial Enumeration without Blocking Clauses

### 18 Dec 2023 v3

@@ Introduction
Notice that the main loop is identical to an AllSAT solver based on non-chronological CDCL; the only differences are embedded in the procedure to get the conflict and the partial assignments. (From now on, we color in red the lines that differ from the baseline CDCL AllSAT solver.)
@@
$\{\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{1},}\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{3}}\},\{{\color[rgb]{0.00,0.00,1.00}\neg x_{1},}\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{3}}\},\{\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{1},}{\color[rgb]{0.00,0.00,1.00}\neg x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{3}}\},$
$\{{\color[rgb]{0.00,0.00,1.00}\neg x_{1},}{\color[rgb]{0.00,0.00,1.00}\neg x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,1.00}x_{3}}\},\{\phantom{\neg}{\color[rgb]{1.00,0.00,0.00}x_{1},}\phantom{\neg}{\color[rgb]{1.00,0.00,0.00}x_{2},}{\color[rgb]{1.00,0.00,0.00}\neg x_{3}}\},\{{\color[rgb]{1.00,0.00,0.00}\neg x_{1},}\phantom{\neg}{\color[rgb]{1.00,0.00,0.00}x_{2},}{\color[rgb]{1.00,0.00,0.00}\neg x_{3}}\},$
$\{\phantom{\neg}{\color[rgb]{0.00,0.39,0.00}x_{1},}{\color[rgb]{0.00,0.39,0.00}\neg x_{2},}{\color[rgb]{0.00,0.39,0.00}\neg x_{3}}\}$
@@
$\{{\color[rgb]{0.00,0.00,1.00}x_{3}}\},\{{\color[rgb]{1.00,0.00,0.00}x_{2},}{\color[rgb]{1.00,0.00,0.00}\neg x_{3}}\},\{{\color[rgb]{0.00,0.39,0.00}x_{1},}{\color[rgb]{0.00,0.39,0.00}\neg x_{2}},{\color[rgb]{0.00,0.39,0.00}\neg x_{3}}\}$

### 09 Jan 2024 v4

@@ Introduction
Notice that the main loop is identical to an AllSAT solver based on non-chronological CDCL; the only differences are embedded in the procedure to get the conflict and the partial assignments. (From now on, we underline the lines that differ from the baseline CDCL AllSAT solver.)
@@
$\{\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{1},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{3}}\},\{{\color[rgb]{0.00,0.00,0.00}\neg x_{1},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{3}}\},\{\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{1},}{\color[rgb]{0.00,0.00,0.00}\neg x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{3}}\},$
$\{{\color[rgb]{0.00,0.00,0.00}\neg x_{1},}{\color[rgb]{0.00,0.00,0.00}\neg x_{2},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{3}}\},\{\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{1},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{2},}{\color[rgb]{0.00,0.00,0.00}\neg x_{3}}\},\{{\color[rgb]{0.00,0.00,0.00}\neg x_{1},}\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{2},}{\color[rgb]{0.00,0.00,0.00}\neg x_{3}}\},$
$\{\phantom{\neg}{\color[rgb]{0.00,0.00,0.00}x_{1},}{\color[rgb]{0.00,0.00,0.00}\neg x_{2},}{\color[rgb]{0.00,0.00,0.00}\neg x_{3}}\}$
@@
$\{{\color[rgb]{0.00,0.00,0.00}x_{3}}\},\{{\color[rgb]{0.00,0.00,0.00}x_{2},}{\color[rgb]{0.00,0.00,0.00}\neg x_{3}}\},\{{\color[rgb]{0.00,0.00,0.00}x_{1},}{\color[rgb]{0.00,0.00,0.00}\neg x_{2}},{\color[rgb]{0.00,0.00,0.00}\neg x_{3}}\}$